Oxygen gas having a volume of 1080 cm3 at 22.5°C and 1.01 x 105 Pa expands until its volume is 1550 cm3 and its pressure is 1.10 x 105 Pa. Find (a) the number of moles of oxygen present and (b) the final temperature of the sample.
Answer #1

(a)

Ideal gas law:

pV= nRT

Here,

n= number of moles

p = absolute pressure = 1.01× 105 pa

V= volume = 1080 cm3= 0.001080 m3

R = Gas constant = 8.314472 J/mol-K

T = absolute temperature in K = 22.5+273 = 295.5 K

n = PV/RT= (1.01× 105)(0.001080)/(8.314472)(295.5)

n = 0.044 moles

(b)

Expansion from state 1 to state 2

p1V1/T1 =p2V2/T2

T2= p2V2T1/p1V1

= (1.10×105)(0.001550)(295.5)/ (1.01× 105)(0.001080)

= 461.88 K = 188.88 °c